FEL Monthly Highlights
May 2001

Excellent progress was made this month on diagnostic instrumentation and the dipole magnet design efforts for the FEL upgrade. The contract for the corrector dipole magnets has been placed. We received the upgraded power supply for the electron gun high voltage system.

Our current analysis (by Dave Douglas) of the energy acceptance of the electron beam transport system in the recirculator is looking very good compared to the nominal design value of 10%. See the WBS 3 report for details.
We began the process of recertifying the Laser Safety System (LSS) for the start-up of the summer operations period.

We thank Eli Zimet for releasing an additional $250k of ONR FY01 funding for the Phase 2 upgrade effort.  This brings the project budget to within $50k of the original budget ($4.5M) of the phase 2 work statement.  This document will be adjusted accordingly.  The project team continues to work with the JTO office to firm up the
contract and start date for the JTO laser materials studies which we hope begins this month.  We are also working with JTO and JTO’s contractor on the technical planning for the high power FEL workshop scheduled for June 5-6.

Abstracts for the FEL conference were prepared and reviewed prior to submission by the deadline of 5/11/01.

We hosted a visit to the FEL, the cavity fabrication facility, and the optical klystron test area by Cdr. Roger McGinnis, Neal Barron, Quenton Salter, and two naval architects to determine the feasibility of adapting our design to a shipboard configuration.

We also hosted three visitors from DARPA (Tim Grayson, Gail Heim and Sheldon Meth) for a tour and briefing on the FEL program.

Project Cost Performance:
The project (Phase 1) budget for the period June 1, 2000 to September 30, 2001 is $9,029k.  The Phase 2 project was approved for $4,500k for a performance period of February 1, 2001 to September 30, 2002.  The project through the month of May has a total of $3,934k of performance scheduled (assuming the project started at the
originally planned start date of April 1, 2000).  The work performed through the end of May was $4,427k, which is 49% complete vs 44% scheduled.  The actual cost accrued through April totals $5,497k.  This results in a schedule variance of +$493k and a cost variance of -$1,069k. We are on track for having the first phase of the project
fully obligated by September.  (The total project cost to date-actual costs plus commitments is $7,141k).  A negative cost variance of ~$1M is presently showing but this is largely an artifact of having to obligate funds early in the project in order to ensure full obligation and costing by the project end date.

WBS 3 (Beam Physics):
Correction of a misperception in the DIMAD analysis of recirculator performance has suggested the driver may accept a FEL-induced momentum spread in excess of the 10% design specification.  As a result, the system may allow operation at FEL extraction efficiencies well beyond the nominal 1% value, with an associated increase in output
power.  Estimates characterizing pi-bend specifications were documented in JLAB-TN-01-024.  A review of the QX quadrupole systematic multipole specification was conducted.  As the bore is larger than that in the Demo, the specification can be relaxed from 1 ppt to 3 ppt without untoward performance effects.  The prototype magnet meets the revised specification. The review is documented in JLAB-TN-01-027, available on line at:

WBS 4 (Injector):
Completed drawings for the shield assembly (door swing mechanism) and submitted to the shop for manufacturing. At month’s end these parts were fabricated and awaited welding. Signed off the remaining shield door test drawings. The ceramic stack pieces have been fitted and stacked, awaiting welding.  Design continued on the interface between the gun and light box, defining an internally tapered tube at the exit of the gun.  Continued detail efforts on the cathode support tube and stalk. The gun vacuum chamber drawings were reviewed and returned to the designer for revisions.  The drawings are going to a manufacturing sequence series.  The anode plate is being polished to 1 um.  Received the 6" RF shielded gate valve from VAT. Submitted the 10 to 6 MV/m spacer to the shop for fabrication.

Installed a 1um stainless steel faraday shielded implanted electrode into the Field Emission Test system. Ran voltage on the 1um SST implanted electrode (implanted with a faraday shield in place).  Preliminary results show much higher emission currents than the electrodes implanted without the faraday shield

We received the gun high voltage power supply from Glassman ,Inc by the end of the month.

WBS 5 (SRF):
Vendor site visit to Aboloty Engineering to check on the vacuum vessel and space frame fabrications.  Space frame was found out of tolerance and vendor is making plans for straightening vacuum vessel top, but flanges were found out of tolerance and vendor is making plans for a fix.  Vendor site visit to ADC, thermal shield vendor.  Inspected
parts, material, QA program, and project planning and integration with other contracts.  All fabricated parts, materials and programs are in good order.

Miscellaneous cryogenic piping package complete and the SOW has been signed off. Received and evaluating quotes for tuner fabrication.

The weld development for the HOM couplers are complete with the first two prototypes completed.  Cavity tests have shown a sloping Qo value on cavity JL005 with full chemistry.  This problem has lead to several tests to determine the cause of the slope.  So far tests have indicated that the problem is not the indium seal and the
next test is to heat-treat cavity JL005 to 800C for 6 hours to determine if niobium hydride precipitation (Qo disease) is the problem.  Flange leaks have developed on the FPC flange and testing to resolve the issue has started.

WBS 6 (RF):
Zone 1 – Operational.
Both HVPS's were cleaned inside of any dust and debris of prevent random arcing.  The brushes for unit 4 were replaced during the last run.  Any bad brushes in unit 3 will be changed next week.
The brushes for unit 3 were replaced this week.  Testing of Zone 1 automatic ramp down voltage not done, due to no resolution on how to lock up the tunnel.

Zone 2 – Operational with no problems.

Zone 3 – The circulators were tested this month.

Zone 4 – RF Control Modules were installed temporarily for an experiment to test the stability of the klystrons by adjusting the filament voltage to control klystron beam current.  The missing circulator for position 8 will be installed as time is available.  This week, 7 of the klystrons were operated to determine the stability of their cathode current and hence RF power when the heater voltage is set much less than normal.  The results are not yet available.

Gun HVPS – The vendor (Glassman) shipped the upgraded HVPS by the end of May.  The design of the
HVPS Tank is in progress.  A new cylindrical tank is being considered as a replacement for the elongated tank design.  The design of a new elbow for the transmission line is just getting started.

Quarter HVPS – PSC has submitted a quote of $279K total to upgrade the 2 HVPS's for the Quarter.
After talking with our internal engineering group (HPEE), we decided to build this Power supply in house.
Injector Klystrons - In performing a helium leak test on the collector of the first 100kW klystron, CPI, feels they may have received parts that were not made of the specified oxygen-free copper.  They are making additional tooling to make a definitive test.  This second test is delaying the project 3 weeks.  If they have to remake the
parts with new copper, the project will slip 3 to 4 months according to their first estimates.

WBS 8 (I&C):
The drive laser pulse truncator specification is complete and the design is proceeding. Many thanks to Curt Hovater, Trent Allison and the EID group for their design assistance.  This device will insert between the existing Drive Laser Pulse Controller (DLPC) and allow for electron beam macropulses (in 500 nanosecond steps) between the existing 1 microsecond and 10 microsecond selections that currently exist.  This will also allow for macropulse frequencies to extend beyond the existing 60 Hz to 100 Hz, 300 Hz, 1 kHz, 3 kHz, 10 kHz, and to 20 kHz. During the original FEL design it was doubted that the FEL could even lase in pulsed mode, hence the gap between 1 and 10 microseconds (go figure!). This will be inserted between the DLPC and the ConOptics 305 mode locker controller. There will be no compromise to the MPS since it can only "pull down" the E/O cells and not supply power to gate them ON. This truncator will operate in pulsed, CW and in Alignment mode. The important feature is that it will be synchronous with the FEL output pulses; that is phase stable for experiments like Bob Austin's who require the 18MHz signal to gate the cryodetectors for enhanced noise immunity. As always Al Grippo will create a simple EPICS screen so all can enjoy.

The viewports have been ordered for the upgrade. An order for 40 pcs.; 4.62 flange fused quartz viewports with a clear aperture of 2.69" was placed.  Once these arrive they will have 4 "blind tapped" holes added for mounting the beam viewer camera assembly, then have a gold coating sputtered on. The shop will most likely do the holes and Larry Phillips will supply the Au.

A Golay Cell was also ordered, this will function as a spare for the Happek interferometers and also be used to evaluate THz radiation from a port at the 2F region. A diamond window from ALS was installed at 2F03; the exit of the upstream optical chicane. This port will be used to evaluate the THz radiation emitted from the chicane bend.  The X-Ray detector's, downstream of the wiggler, Pb shield was installed and the control/software was updated (moved over from the scrapers) -the "X" (up/down) & "Y" (right/left) have been adapted to suit the newly reincarnated
operational modes required from the backscattered X-Ray experiments.  The DSP56F805 microcontroller software is proceeding well.  This device is slated to be used for a number of applications; BPMs, OBPMs, trim PS cards, and for some simple motor control. The Code Warrior software, including compiler and debugging tools have been installed. A sample program has been edited, compiled and loaded with and observed running in the DSP, with variables in memory changing in accordance with the program.  A simple program was also coded to read a built-in A/D channel.  Program appears to work, but have not yet applied a voltage to see the data change.

Progress has been made for remote control of the mirror test stand YAG laser; the initial version of EPICS software for monitor/control is complete and being debugged.  The coding for the (new) VME interface board device driver is complete.  The AVME9510 has all of the control and status functions in a single board (& at a reasonable $);
digital in/out and analog in/out.  The field wiring design is being worked out and should be installed next week.  The LSS connections have been made and tested.  The X-Ray DC Motor has been wired up and is operational through EPICS.  The software control is also completed and being tested for 2 DC motors to be used to control x-ray
detector position.

The Lab #2 LSS low level certification passed (although it was only done for training purposes). There exist NO known problems with the LSS master, or any of the labs (except #5 which has not been commissioned).  The new hutch in lab #6 is complete and awaiting a decision on how and where the penetrations should go. Test boxes were built for testing the Beam Viewer Chassis and the PLC in the LSS system.  The next effort on the LSS is to clean up the overall system print.

There has been more improvements to the website http://laser.jlab.org to include a feedback link, technical document links and more system information.  Each of the systems is getting an overhaul - this is critical to have an accurate "as-built" reflecting all of the changes made over the last couple years prior to the "rip-out".

The Molectron quad detector was tested; it performed very well.  This will be used for optical transport control. EPICS software was written for a CNC-like script which a user can use to program the FEL to perform a fairly general set of drilling motions.  This has not yet been tested

The last of the LSS work is being finished up.  New status signs were put on the doors of labs 5 and 6. The new hutch in lab 6 had limit switches put on the doors. There are 12 doors, each has two limit switches - one for each corner opposite the hinge.

A productive visit was made to Triumf regarding their CAN bus implementation.  They have ~ 500 devices installed and operational on the Controller Area Network (CAN). It is our intention to use this for the new BPM electronics and for the power supply control for the upgrade.  The new VME CAN interface also arrived this month, it is the
same one that is used at Triumf (and supported in EPICS).  All major components are on site for this BPM prototype, the '807 DSP (samples just released from production this week), the Philips CAN interface (to the DSP), the RF filters, and the Analog devices RF-to-DC converter. The only thing that remains is the 3/3.3 volt power
management/conversion.  These will be tested during the summer run. The Picomotor Mux Chassis has been re-installed in the control room.

Several drawings are in EECAD:  Ion Pump Power Supply Main Board Schematic; Molectron OBPM and daughter board Schematic; 64 Bit General purpose Digital I/O card; Single channel OBPM Amplifier Schematic; Picomotor Relay and Ion Pump Power Supply Wiring Diagrams and the Analog Differential Driver Board Schematic as well as several pages of the MPS System Drawing.  The following drawings have been received for review/signature: OBPM enclosure box; Upstairs Dump Raster Board Schematic; Picomotor Relay Chassis Mechanical; and the LSS System Block Diagram.

Optical BPMs are making great progress. All testing has been done using the drive laser and turning mirrors on picomotors.  The Auto zero function works very well; this is used once the optical transport has been setup the detector is moved to zero rather than moving the optical transport to the same position for each lasing setup. The
transport lock chassis has been checked out, the last remaining thing is the Digital-to-Analog (D/A) converter. This has real time control over six axis of transport mirrors (X & Y for Coll., can 2 & 6).

WBS 9 (Transport):
Injector Dipoles (DU/DV):
o  DULY Research, using small corrections to their magnetic model matched the core field and integrals of both the
    small and large Injector Dipoles at the same current, all within the tight 1 part in 10,000 of the required by Dave
    Douglas’ specifications.
o Using their magnetic model, they probed the sensitivity of their sliding shorts to the field clamps solution to uniform
    field integral of the small injector dipole. They found that changes of a mm keep the integral within specification.
o They sent the complete package of drawings both the Large (DU) and the Small injector Dipole (DV) to us for
Optical Chicane Dipole (DW):
o  The design package is finished backchecking for the second time.
o  Innovation Services reviewed our coil specification for dipole insulation, bringing it up to date for resin systems and
    enhancing the quality control provisions.  They are now working on a final report that will serve as an in house
    reference on how they arrived at their conclusions embodied in the revised specification.
o We received the electrical steel pole tip material for this magnet and the 108° Dipole from the custom specified
    roll that had the best thickness uniformity characteristic.
Arc Dipoles (GY, GX, GQ):
o  The magnetic analysis of the GY with path length correctors and now synchrotron light periscopes continued.  AES
    was successful in reducing the local field perturbation caused buy the large 1.5 inch diameter holes for the
    diagnostic periscopes.  The electrical steel was extended around the hole.  AES was also successful in widening the
    good field region to within specification by changing the inner edge of the electrical steel.  AES decreased the size of
    the path length correctors.  The smaller required slots reduced the perturbation to the field so the horizontal parallel
    offset of the beam reduces to sub mm values. This is probably the final solution for this magnet.
o  AES started magnetic modeling of the Arc Bend Dipole (GX). They removed the nubbins on the edges of the
    poles and obtained the 1 part in 10,000 bulk field and integral uniformity required of the specifications. However
    they have to work on raising the bulk field.  They also activated just the two of the four coils that face the gap in
    order to investigate the half-field switch-to-UV configuration.  With this they were also successful in obtaining the
    appropriate field uniformities.  This success allowed continued design of the current leads for the coils to enable this
    form of powering.
o  Analysis of the results of the magnetic model of the Arc Bend Dipole (GX) showed incorrect field quality in the
    bulk field, longitudinally.  Their results showed a parabolic change in the bulk field of the magnet, lowering by 7
    parts in 10000 over the 0.6 meter central zone.  At month’s end, consultations with David Douglas and
    further efforts are on going to find out if this is acceptable, to see if we experienced similar results with the IR Demo
    and to determine what Bulk Field really means when applied to a real magnet.
o At the same time, the assembly drawing and detail drawings were firmed up. The coils were modeled and the
    field clamps were developed. As the month closed, design efforts concentrated on introducing the lead configuration
    that allows powering the two sets of coil pairs independently.
o We developed the bus bar mounting scheme to allow upper and lower halves of the coils to be powered
    independently and allowed the magnets to be flipped over 180° with the same hardware.  Because the magnet is
    an asymmetrical four sided figure; it is used right side up in two locations and flipped over in the remaining two
    locations.  Our design makes a universal magnet that can be used in both orientations.
QGh Procurement on how best of purch (3 inch quad):
o  The coil manufacturing bids were received and the evaluation team given the papers to start their process.
o  The prototype quadrupole had cycles of chamfer cutting to reduce the n = 6 pole, our final requirement of
    the prototype process.  The results indicate a simple 0.187 inch chamfer will almost meet David Douglas’s
    specification.  Additional simple chamfer appears to make the n =10 pole rise.
o We added end chamfer shim to the prototype quadrupole.  It showed promise in understanding the reduction of the
     n = 10 pole and another run with double the shim showed double the results.  We will now try a series of cuts using
    this understanding to reduce both the n = 6 pole and the n = 10 pole.  David Douglas has found that he can live
    with the higher 3 part per thousand errors of the original configuration but is encouraging us to get a lower value.
o The core drawing and statement of work were signed as the procurement package was sent out for bid.
QX (Trim Quad):
o The 3D magnetic model was debugged and is now yielding good results.  Now detailed magnetic design can start,
    moving the protrusion of the coil ends in the horizontal or vertical to minimize the higher order multipoles.
o We reduced the length of the magnetic model by one inch and obtained a better overall quadrupole gradient
    integral.  The reduction was indicated to back the coils away from the field clamps of the GQ dipole.
o We received the conductor that was specially ordered for these coils.  It was slightly out of our expected tight
    tolerances but may now be sequestered and the design based on it, which will significantly reduce the variability of
    the final magnets.
o We are integrating the sequestered conductor dimension into our 3D magnetic model and into the design drawings.
    We are working on a plan to individually match coil size (we have only 8 magnets to build) that results from a 140
    conductor build up into a variable magnet geometry that will yield the best magnetic characteristics.
Corrector Dipoles (DB & DJ):
o We placed the contract with Milhous Co. of Lynchburg VA after they clarified their manufacturing plan.
o  Robin Wines returned from leave to take up design of the sextupole, expanding from her successful 2D design into
    a 3D design in TOSCA.
o The lower half of the new alignment cartridges were put on order.
o  A single Dipole Steel bid was received from a reputable company.  The folks in procurement are following up with
    contacts with the other vendors to get comments.
o We are anticipating a response today from the company so that we can place that procurement for steel.
o We continued design layout of the girders between cryomodule.
o We completed design layout of the girders between cryomodule
o The Task definition to be used by an engineering service contractor for design.
o The heavy-duty alignment cartridges from CERN (European Organization for Nuclear Research) for use under the
    large 180 degree dipoles are being shipped.
o We are now assembling the attachments for the task definition to be used by an engineering service contractor for
    design of the arc chambers.
o Jeff Karn gave us notice that he will be leaving the Lab on the 20th of June.  Robin Wines will take over direction of
    the magnet effort and Tom Hiatt will take over direction of Magnet Measurements.

WBS 10 (Wiggler):
Previous magnet measurements showed the field integral in the dispersion section to be non-zero.  A temporary trim coil was wrapped around all poles to provide a "tweakable" offset. Preliminary measurements show such a coil can be used to zero out the integral.  Additional measurements are in the works.  In the wiggler, temporary coils were configured on the field clamps to compensate for the "missing pole."  Magnet measurements showed the extra trim coils could be tweaked to balance pole fields.  G. Biallas is pursuing building production coils.  C. Hutton has resumed the design of the test stand hardware.

Fabrication of the corrections coils for the wigglers has started.  The design of the coils is being incorporated into the overall design.  Most of the drawings for the wiggler measurement jig have been signed off.  Characterization of the dispersion section with a corrector coil continues.

Measurements of the dispersion section were made to characterize the linear behavior of the trim coils.  Data is being reviewed.  We are still waiting for the production trim coils for the wigglers.  Drawings for probe mechanical drive were signed off and parts are out for order.  Still in the works is the drawing for the Hall Probe mount.

WBS 11 (Optics):
The design for the ultraviewer is complete; we anticipate the drawing package going out for bid early next week.  Prints for new, near Brewster windows for the User Labs have been signed and are going to Procurement.  The fabricator for the prototype deformable mirrors has proposed several changes that will result in cost savings and
accelerated delivery.  (What more can a WBS manager ask for?).  The rest of the internal mounting components for the mirror test stand arrived and are being cleaned.  The manufacturer of our laser interferometer, Veeco, has completed the metrology of  the 3" diameter silicon mirrors produced by RMI. They found that the mirrors met or
exceeded specifications, qualifying RMI as a source for substrates for the IR Upgrade.  It was also an excellent test of the interferometer, because some of these measurements were near the limits of its capabilities. This gives me confidence it will perform well as part of the Mirror Test Stand. We set up a pickoff in the Drive Laser Enclosure that allows I & C to test OBPM detectors.  It will also permit testing of the automatic stabilization features of the FEL laser beam before we actually use it.

The ultraviewer and near-Brewster window drawing packages are out for bid, we are already ordering associated optics for the assemblies.  The design focus shifts back to the optical cavity assembly internals, in particular the gimbal mounts.  We have generated more detailed specs on the linear motion needed to orient the mirrors, and have started detailing the linear actuation.

The design on the optical cavity assembly internals is proceeding For mirror motion, we have drawings on the actuation, but will also pursue a scheme hereby we use a commercially-available motorized actuator.  Several companies produce these, but none will rate their units to a pressure of 10E-8 Torr.  We will borrow one and do base
pressure and outgassing tests during the next several weeks.  The Nd:YAG laser for the mirror test stand was tested this month.  It made specified power easily, and the interlocks tested OK.  The beampath shields arrived, and one was installed, another is in the shop for modification.  Motion control of the gimbal mount was installed and tested.  Parts to hold our 3" DFM assembly in the 6"gimbal mount were cleaned for HV use and are ready to be used.  We began building a scaled-down version of the mirror metrology system.

Chris Behre, an undergraduate student from ODU, joined the Optics Group for the summer. The design on the optical cavity assembly internals continues.  We worked on better defining the actual components for the linear actuation.  We held discussions with the makers of lvdts and motorized actuators.  We should receive test articles of both components next week, and will do outgassing studies.  The rest of the beampath shields for the mirror test stand arrived, and we will complete installation after we finish some items required to begin Operations.  We received the modified IR Demo mirror mount, so we can better quantify a number of parameters associated with our edge-cooled scheme.  These will be needed for the outcoupler mirror on the Upgrade.  The optical cavity assembly design has progressed in the following areas:  The movement of the gimbal mounts has been converted from linear translation stages to precision rails and ball bushings.  We decided to reduce some complication in the framing by attaching the rails to the side of the vacuum enclosure.  The gimbal assembly support structure, which also supports the linear actuators and lvdts, has been designed.  We received quotes and engineering details on another UHV motorized
actuator.  The actuator surpasses our specifications, and should integrate easily into our design.  We are checking with at least two other vendors to confirm the pricing is competitive.  We received a motorized actuator and will start outgassing studies next week.  We received refigured, broadband (5.0 - 6.6 um) outcouplers, and are inspecting

I attended a Joint Technical Office Technical Area Working Group, held in Albuquerque the 23rd and 24th. While there, I gave presentations on laser/materials interactions work done here, and on the User Facility itself. Attendees were quite interested in the unique combination of high average power and ultrashort pulsewidth.  The workshop
also gave me an opportunity to discuss support by staff from NAVSEA (Dahlgren) and NRL for the series of experiments we'll do under the JTO contract.

Work continued to define the "killer" demonstration of our high flux X-ray source.  We believe a pump/probe experiment looking at resonant phonon decay in crystals might be the answer.  We are continuing to analyze the requirements. We prepared hardware for installation of:  a)  a diamond window to allow extraction of far IR/mm waves, b)  a new X-ray diffraction crystal, and c)  silicon viewers before and after the first 180 bend to allow careful CSR emittance growth measurements to be performed at a later date.  No operations early in the month.  However, we did have put to put a temporary work procedure in place for working in the FEL Building today and yesterday after a calibration problem was discovered with the oxygen sensors that are used for ODH detectors wherever cryogenic gases are being used.  The calibration problem was noticed during a scheduled spill test in the CEBAF accelerator.  Re-calibrations of the detectors and re-certification of the ODH system are expected by next week.  Many groups in the Accelerator Division worked hard and expeditiously to solve this problem.  Thanks to all involved.
Preparations continued for the summer user run including check-out of diagnostics and LSS certification.  We also continued working with various groups in the Lab to help with the re-certification of the ODH sensors. The Safety Systems group has identified a recalibration procedure that will allow the existing ODH sensors to be recalibrated
for the 18% oxygen concentration set-point.  This procedure was successful.  The following work was done this week to prepare for operations for the summer run:  Installed insertable mirrors in U.L. 3 to divert beam from the metals processing hutch to the nanotubes experimental setup.  Installed manual beam stops in U.L. 3.  Aligned
FEL OTS.  Installed insertable mirror in U.L. 2 beam dump assembly. Installed OD 1 filters in OBPMs (FL02 and FL08).  Aligned optics on OBPM test stand in the Drive Laser Enclosure.

Privacy and Security Notice